OpenStudy (anonymous):

What is the equation in slope-intercept form of the line through the points (2, -3) and (-3, 7)?

OpenStudy (anonymous):

y-2=(7+3)/(-3+2)*(x+3) y-2=-10(x+3) y=-10x-28